So I've been given the following values from an experiment where we ran water through a pipe and were told to observe only the Q(gpm) and the change in Pressure(psi). From there, we need to calculate the head loss, s-value, f-value, and the Hazen-Williams coefficient. So I need help figuring out which equations to use for those?
Answer #1

Ans) According to question , flow rate (Q) and change in pressure (\DeltaP) are known and we need to determine head loss , slope Darcy friction factor (f) and Hazen-William coefficient (C) .

Apply Bernoulli equation between between any two point inside pipe ,

P1 / \gamma + V12 / 2g + Z1 = P2 / \gamma + V22 / 2g + Z2 + HL

=> HL = \Delta P/\gamma + (V12 - V22) / 2g + \Delta Z

Velocity can be determined by using Q = A x V , where A is cross sectional area of pipe = (\pi/4) D2

From above equation , head loss can be determined

Also, using equation HL = f L V2 / ( 2 g D) , we can determine Darcy friction factor (f) = 2 g D HL / L V2

Also, HL = 10.67 L Q1.85 / C1.85 D4.87

where, C = Hazen William coefficient (C)

Using above equation, we can determine Hazen William coefficient (C) ,

Now, slope (s) can be determined using below equation,

V = k C R0.63 s0.54

where, k = conversion factor = 1.318 for US unit and 0.849 for SI unit

R = Hydraulic depth = D/ 4 ,where D is diameter of pipe